# Permutation invariant matrix quantum thermodynamics

### v1

@@ 1 Introduction
An important strand of research in gauge-gravity duality investigates the relation between the gravitational thermodynamics in AdS and the CFT thermodynamics. The transition in semi-classical gravity between the vacuum AdS solution and the AdS black hole geometry was shown to imply a phase transition at large ’t Hooft coupling in the large $N$ limit of $N=4$ SYM theory [4]. Motivated by a study of phase structures in weakly coupled gauge theories, it was shown in [8] that the $U(N)$ invariant sector of the 2-matrix harmonic oscillator has an exponential growth of the degeneracies at low energy levels and a finite Hagedorn temperature in the large $N$ limit. This mirrors properties of weakly coupled gauge theories [9] in the large $N$ limit. Finite $N$ partition functions for 2-matrix quantum mechanics were analysed for $N$ up to $7$ [10] and the Hagedorn transition related to a convergence of zeroes to a point on the real axis $x=e^{-1/T}=e^{-\beta}$. Recent work has shown that super-symmetric matrix models with known gravity duals can have non-supersymmetric analogs which capture similar large $N$ thermodynamics [11, 12].
@@
The study of permutation groups as manifest, as opposed to hidden, symmetries in matrix models has been initiated in the context of applications of matrix models to ensembles of matrices in computational linguistics [27]. The general 13-parameter Gaussian model [28] has been shown to describe approximate gaussianity in matrix representations of words [29, 30] and a reduction to a 4-parameter model appropriate for symmetric traceless matrices has been used to demonstrate approximate gaussianity for statistical correlation matrices [31].
@@
In [6] we gave, following earlier work in the context of the BFSS model [13], the path integral formulation for the partition function of permutation invariant matrix models, by discretising the Euclidean time direction so that parallel transport is implemented by group elements as in lattice gauge theory, and taking a continuum limit. This allowed us to recover the Molien-Weyl formula in the Gaussian case. In [7] we derived explicit formulae for the partition functions of the general 11-parameter Gaussian permutation invariant harmonic oscillator quantum mechanical systems.
@@
The basic mechanism responsible for the region of negative specific heat capacity, is super-exponential, or as we will find, near-exponential growth of degeneracies at low energies $k\lesssim N\log N$. At higher energies, finite $N$ effects tame the rapid growth of degeneracies and the micro-canonical specific heat is positive.
@@
  • We find indications of similar thermodynamic characteristics to the gauged permutation invariant matrix model in a system with gauged $U(N)^{\times 3}$ acting on complex $3$-index tensors. This includes a Hagedorn-like cross-over transition which is expected to vanish as $N$ goes to infinity, in-equivalence between the canonical and micro-canonical ensembles in the low temperature region, and negative specific heat capacity at low micro-canonical temperature. The evidence involves a combination of the high temperature behaviour we have derived from the path integral formulation, and computational evidence based on known group-theoretic formulae involving Kronecker coefficients (see eqn (8.7)) for the finite $N$ counting of tensor model invariants. The SAGE code needed to do the computations is included in the discussion.
@@
Section 6 discusses the path integral formulation of vector and tensor models and we exhibit sample small $N$ partition functions. Subsection 6.5 deals with the charge-zero sector of two matrix models. We derive a Molien-Weyl formula for the partition function and evaluate explicit expressions for partition functions for $N=2,3$ and $4$ and the large $N$ low temperature limit observing that in the micro-canonical ensemble this charge-zero system also exhibits negative specific heat (see section 9.2). Section 7 deals with the high temperature scaling in the path integral formulation.
In section 8 we combine information from the high temperature limits with representation theoretic formulae in two cases of interest, to deduce that there is a negative specific heat capacity at low temperature in the micro-canonical ensemble followed by a transition and a high temperature phase of free oscillators. The two cases of primary interest involve a 3-index tensor system and a complex-matrix system with zero charge.
Section 9 presents a discussion of general forms of degeneracies leading to negative specific heat capacities along with connections with the physics of small black holes via the AdS/CFT correspondence. This discussion includes instances of multi-matrix thermodynamics which show negative specific heats when the number of matrices scales with the energy. We conclude with a summary of our results and a brief discussion of future research directions in section 10. The Appendices develop technical points arising in the main discussion.
@@ 3.1 Super-exponential degeneracies and vanishing Hagedorn temperature
diverges for any finite $x$ or $\beta$, or equivalently any finite temperature. There is a similar growth of degeneracies in the stable region in the context of tensor models, and this has been interpreted as a zero-temperature Hagedorn transition [41] [42]. We will return to the tensor model case in more detail in sections 6 and 7.
@@ 3.2 Thermodynamic quantities in the canonical ensemble
$\displaystyle{\cal C}_{{\rm{sh}}}={\partial{\cal U}\over\partial T}$ (3.9)
@@
$\displaystyle C_{{\rm{sh}}}={\partial U\over\partial T}={1\over N^{2}}{\cal C}_{{\rm{sh}}}$ (3.10)
@@
The probability of energy levels $E$ at temperature $T$ in canonical ensemble is given by
@@
$\displaystyle{\cal S}(T)$ $=\langle\log P(E)\rangle={(E-{\cal F})\over T}=\log{\cal Z}+{{\cal U}\over T}$ (3.22)
@@ 3.4 Zeroes in the complex plane
4 PIMQ-Thermodynamics in the micro-canonical ensemble : Negative specific heat and (in)equivalence of ensembles
@@ 4.2.1 Energy versus micro-canonical temperature
Conjecture 3: It is reasonable to conjecture that in the large $N$ limit, the value of the un-normalised micro-canonical energy at the critical temperature where the specific heat capacity diverges, i,e $k_{\rm{crit}}$, obeys
@@ 4.2.2 Specific heat capacity versus micro-canonical temperature
$\displaystyle C=-{(\Delta\log Z(k))^{2}\over\Delta(\Delta\log Z(k))}$ (4.15)
@@ 4.3 Equivalence and in-equivalence of ensembles
Expressing $E$ as a function of $T_{\rm{micro}}$, we can obtain a plot of $S_{\rm{micro}}(T_{\rm{micro}})$. As discussed around equations (3.31), we can also obtain an entropy function $S(T)$ for the canonical ensemble. Superposing $S_{\rm{micro}}(T)$. Superposing these plots in Figure 16 we obtain the expected feature discussed above of ensemble in-equivalence at low temperature followed by equivalence at high temperature.
@@ 5 High temperature expansion of PIMQ-Thermo and small cycle dominance
In this section we will analyse the behaviour of the canonical partition function ${\cal Z}(N,x=e^{-\beta}=e^{-1/T})$ in the infinite temperature limit where $x\rightarrow 1$. The behaviour in this limit is controlled by the singularity as $x\rightarrow 1$. Given that ${\cal Z}(N,x)$ is a sum over partitions $p$ of ${\cal Z}(N,p,x)$ which, as we show, takes the form
@@
where ${\rm{Deg}}(N,p)$ is an integer we will define and ${\cal R}(x)$ is a regular as $x\rightarrow 1$. We will show that the leading singularity of ${\cal Z}(N,x)$ as $x\rightarrow 1$ comes from $p=[1^{N}]$. The next-to-leading singularity come from $[1^{N-2},2]$, followed by other cycle structures involving a large number of $1$-cycles and a few small cycles. The proof that $p=[1^{N}]$ gives the leading singularity follows, while the proof that the next-to-leading singularity comes from $p=[2,1^{N-2}]$ is given in Appendix B.
@@
This means that where we have separated the terms singular at $x=1$
@@ 5.1 Cycle structures and the high temperature expansion
For $k=1$,we have ${\rm{Deg}}(N,[2,1^{N-2}])=N^{2}-2N+2$. Note that the coefficient of $N$ is $-2B([2^{k},1^{N-2k}])$ where $B$ is the branching number of the partition $[[2^{k},1^{N-2k}]$. The branching number of a general partition $p=[\{a_{i}^{p_{i}}\}]$ is $\sum_{i}p_{i}(a_{i}-1)$. It has an interpretation in terms of branched covers and plays a role in the Riemann-Hurwitz formula for the genus of covering surfaces in two dimensions. We will find shortly that this property, that the coefficient of $N$ in the degree is the branching number of $p$, holds for general $p$.
@@ 6 Path Integral for Tensor Models
where ${\cal D}_{\tau}=\partial_{\tau}-iA$ is the covariant derivative with the gauge field $A(\tau)=\sum_{r=1}^{s}A^{r}(\tau)$ being a Hermitian matrix in the Lie algebra of $U(N)^{s}$. A subtlety associated with this system is that though the gauge group is $U(N)^{s}$ only $SU(N)^{s}\otimes U(1)$ has a non-trivial action on the tensors i.e. there is only a single $U(1)$ gauge gield. The remaining $U(1)^{s-1}$ decouple completely from the system. The reason is rather straightforward: $U(N)$ can always be factored into $SU(N)$ and an overall $U(1)$ phase. But the $s$ overall phases act as a single phase, yielding a single diagonal $U(1)$ action. The $U(1)$ gauge fields $A_{k}$ are proportional to the identity matrix and therefore do not rotate the tensor’s index, therefore they combine into with their sum $A_{1}+\cdots+A_{s}$ which is the effective $U(1)$ gauge field.
@@ 6.1 Complex-Vectors
where $P^{(1)}_{1,2d-2}(x;d)$ is a palindromic even polynomial of degree $22d-$ 2, e.g. $P^{(1)}_{1,2}(x;3)=1+x^{2}$, $P^{(1)}_{1,4}(x;4)=1+4x^{2}+x^{4}$ and $P^{(1)}_{1,6}(x;3)=1+9^{2}+9x^{4}+x^{6}$.
@@ 6.2 Two-Tensors
One recognoises that the invariants are built from traces of $X_{i}^{j}=\Psi_{ik}\overline{\Psi}^{kj}$ or equivalently the eigenvalues of $X_{i}^{j}$ whose square roots are the singular values of $\Psi_{ij}$ and the problem reduces to counting the invariants of a single matrix. So the invariants are the singular values of the 2-tensor, which can be thought of as a generic complex two matrix.
@@
6.3 Three-Tensors
@@ 9 Negative SHCs in matrix quantum mechanical models and AdS/CFT
$\displaystyle{\partial\beta_{\rm{micro}}(k)\over k}=-{a\over k^{2}}$ (9.17)
@@ 9.3 Negative SHC with super-exponential growth of degeneracies in multi-matrix models
$\displaystyle Z(n,d,N)=\sum_{\begin{subarray}{c}R\vdash k\\ l(R)\leq N\end{subarray}}\sum_{\begin{subarray}{c}\Lambda\vdash k\\ l(\lambda)\leq d\end{subarray}}C(R,R,\Lambda){\rm Dim}_{U(d)}\Lambda$ (9.34)
The sum is over Young diagrams $R$ with $k$ boxes, constrained to have no more than $N$ rows, and Young diagrams $\Lambda$ with $k$ boxes and no more than $d$ rows. $C(R,R,\Lambda)$ is the number of trivial representations of $S_{k}$ in the decomposition of the tensor product $R\otimes R\otimes\Lambda$ into irreducible representations of the diagonal $S_{k}$, i.e. the Kronecker coefficient for the triple of Young diagrams $(R,R,\Lambda)$. ${\rm Dim}_{U(d)}\Lambda$ is the dimension of the $U(d)$ representation corresponding to Young diagram $\Lambda$. The formula simplifies in the case $d,N\geq n$ :
$\displaystyle Z(k,d,N)=Z(k,d)=\sum_{l=0}^{k}d^{l}P(k,l)$ (9.35)
@@
$=\sum_{\begin{subarray}{c}R\vdash k\\ l(R)\leq N\end{subarray}}\sum_{\Lambda\vdash k}\sum_{p\vdash k}\sum_{q\vdash k}{1\over{\rm{Sym}}p}{1\over{\rm{Sym}}q}\chi^{R}_{p}\chi^{R}_{p}\sum_{\Lambda}\chi^{\Lambda}_{p}\chi^{\Lambda}_{q}d^{C_{q}}$
We used the expression for the Kronecker coefficient as a sum of characters. In the last line, we have converted the sum over $\sigma$ to a sum over partitions $p$ and the sum over $\tau$ to a sum over partitions $q$. $\chi^{R}_{p}$ is the character of a permutation $\sigma_{p}$ in conjugacy class $p$ for the irreducible representation $\Lambda$. We can do the sum over $\Lambda$ using character orthogonality.
$\sum_{\Lambda}\chi^{\Lambda}(\sigma_{p})\chi^{\Lambda}(\sigma_{q})=\delta_{p,q}~{}~{}{\rm{Sym}}~{}p$ (9.39)
@@
$\boxed{~{}~{}~{}~{}~{}~{}Z(n,d,N)=\sum_{p\vdash k}{d^{C_{p}}\over{\rm{Sym}}~{}p}\sum_{\begin{subarray}{c}R\vdash n\\ l(R)\leq N\end{subarray}}\chi^{R}_{p}\chi^{R}_{p}~{}~{}~{}~{}~{}~{}}$ (9.40)
@@
Another reason for exploring negative SHCs in multi-matrix models with $d$ scaling as $n$ is that novel scalings of $d$ in multi-matrix models are known [79] to reproduce certain aspects of the physics of tensor models, in particular the existence of limits dominated by melonic interactions. We have in section 8 that tensor models display negative SHCs at finite $N$, so it is natural to explore this feature in multi-matrix models with novel scalings.

### v2

@@ 1 Introduction
An important strand of research in gauge-gravity duality investigates the relation between the gravitational thermodynamics in AdS and the CFT thermodynamics. The transition in semi-classical gravity between the vacuum AdS solution and the AdS black hole geometry was shown to imply a phase transition at large ’t Hooft coupling in the large $N$ limit of ${\cal N}=4$ SYM theory [4]. Motivated by a study of phase structures in weakly coupled gauge theories, it was shown in [8] that the $U(N)$ invariant sector of the 2-matrix harmonic oscillator has an exponential growth of the degeneracies at low energy levels and a finite Hagedorn temperature in the large $N$ limit. This mirrors properties of weakly coupled gauge theories [9] in the large $N$ limit. Finite $N$ partition functions for 2-matrix quantum mechanics were analysed for $N$ up to $7$ [10] and the Hagedorn transition related to a convergence of zeroes to a point on the real axis $x=e^{-1/T}=e^{-\beta}$. Recent work has shown that super-symmetric matrix models with known gravity duals can have non-supersymmetric analogs which capture similar large $N$ thermodynamics [11, 12].
@@
The study of permutation groups as manifest, as opposed to hidden, symmetries in matrix models has been initiated in the context of applications of matrix models to ensembles of matrices in computational linguistics [27]. The general 13-parameter Gaussian model [28] has been shown to describe approximate gaussianity in matrix representations of words [29, 30] and a reduction to a 4-parameter model, appropriate for symmetric matrices which are constant along the diagonal, has been used to demonstrate approximate gaussianity for statistical correlation matrices [31].
@@
In [6] we gave, following earlier work in the context of the BFSS model [13], the path integral formulation for the partition function of permutation invariant matrix models, by discretising the Euclidean time direction so that parallel transport is implemented by group elements as in lattice gauge theory, and taking a continuum limit. This allowed us to recover the Molien-Weyl formula for the generating function of group invariants in the Gaussian case. In [7] we derived explicit formulae for the partition functions of the general 11-parameter Gaussian permutation invariant harmonic oscillator quantum mechanical systems.
@@
The basic mechanism responsible for the region of negative specific heat capacity, is super-exponential growth of degeneracies at low energies $k\lesssim N\log N$. At higher energies, finite $N$ effects tame the rapid growth of degeneracies and the micro-canonical specific heat is positive. The discussion in section 9 shows that it is also possible to get negative specific heat capacity with near-exponential, but sub-exponential, degeneracies as a function of energy, the key point being, as is well known in the statistical physics literature, the failure of concavity of the entropy as a function of the energy in the micro-canonical ensemble.
@@
  • We find indications of similar thermodynamic characteristics to the gauged permutation invariant matrix model in a system with gauged $U(N)^{\times 3}$ acting on complex $3$-index tensors. This includes a Hagedorn-like cross-over transition at a temperature which is expected to vanish as $N$ goes to infinity, in-equivalence between the canonical and micro-canonical ensembles in the low temperature region, and negative specific heat capacity at low micro-canonical temperature. The evidence involves a combination of the high temperature behaviour we have derived from the path integral formulation, and computational evidence based on known group-theoretic formulae involving Kronecker coefficients (see eqn (8.7)) for the finite $N$ counting of tensor model invariants. The SAGE code needed to do the computations is included in the discussion.
@@
Section 6 discusses the path integral formulation of vector and tensor models and we exhibit sample small $N$ partition functions. Subsection 6.5 deals with the charge-zero sector of two matrix models. We derive a Molien-Weyl formula for the partition function and evaluate explicit expressions for partition functions for $N=2,3$ and $4$ and the large $N$ low temperature limit observing that in the micro-canonical ensemble this charge-zero system also exhibits negative specific heat capacity (see section 9.2). Section 7 deals with the high temperature scaling in the path integral formulation.
In section 8 we combine information from the high temperature limits with representation theoretic formulae, in the case of a complex 3-index tensor model with $U(N)$ symmetry, to deduce that there is a negative specific heat capacity at low temperature in the micro-canonical ensemble followed by a transition and a high temperature phase of free oscillators.
Section 9 presents a discussion of general forms of degeneracies leading to negative specific heat capacities. This leads to the interesting case of the hermitian-2-matrix sector (equivalently one complex matrix sector) of operators, which exists within ${\cal N}=4$ SYM theory, where we find that there is negative specific heat capacity in a zero-charge subspace. We also describe instances of multi-matrix thermodynamics which show negative specific heats when the number of matrices scales with the energy. We briefly review the discussion of small black holes in AdS which have negative specific heat capacity, and which motivate the search for a convincing model of their physics within CFT duals. We conclude with a summary of our results and a brief discussion of future research directions in section 10. The Appendices develop technical points arising in the main discussion.
@@ 3.1 Super-exponential degeneracies and vanishing Hagedorn temperature
diverges for any finite $x$ or $\beta$, or equivalently any finite temperature. There is a similar growth of degeneracies in the stable region in the context of tensor models, and this has been interpreted as a zero-temperature Hagedorn transition [41] [42]. We will return to the tensor model case in more detail in sections 6, 7 and 8
@@ 3.2 Thermodynamic quantities in the canonical ensemble
$\displaystyle{\cal C}_{\rm{hc}}={\partial{\cal U}\over\partial T}$ (3.9)
@@
$\displaystyle C_{{\rm{sh}}}={\partial U\over\partial T}={1\over N^{2}}{\cal C}_{\rm{hc}}$ (3.10)
@@
The probability of each microstate of energy $E$ at temperature $T$ in the canonical ensemble is given by
@@
$\displaystyle{\cal S}(T)$ $=-\langle\log P(E)\rangle={(\langle E\rangle-{\cal F})\over T}=\log{\cal Z}+{{\cal U}\over T}$ (3.22)
@@ 3.4 Zeroes in the complex plane
4 PIMQ-Thermodynamics in the micro-canonical ensemble : Negative specific heat capacity and (in)equivalence of ensembles
@@ 4.2.1 Energy versus micro-canonical temperature
Conjecture 3: It is reasonable to conjecture that in the large $N$ limit, the value of the un-normalised micro-canonical energy at the critical temperature where the specific heat capacity diverges, i,e. $k_{\rm{crit}}$, obeys
@@ 4.2.2 Specific heat capacity versus micro-canonical temperature
$\displaystyle C_{{\rm{sh}};\rm{micro}}=-{(\Delta\log Z(k))^{2}\over\Delta(\Delta\log Z(k))}$ (4.15)
@@ 4.3 Equivalence and in-equivalence of ensembles
Expressing $E$ as a function of $T_{\rm{micro}}$, we can obtain a plot of $S_{\rm{micro}}(T_{\rm{micro}})$. As discussed around equations (3.31), we can also obtain an entropy function $S(T)$ for the canonical ensemble. Superposing these plots in Figure 16 we obtain the expected feature discussed above of ensemble in-equivalence at low temperature followed by equivalence at high temperature.
@@ 5 High temperature expansion of PIMQ-Thermo and small cycle dominance
In this section we will analyse the behaviour of the canonical partition function ${\cal Z}(N,x=e^{-\beta}=e^{-1/T})$ in the infinite temperature limit where $x\rightarrow 1$. The behaviour in this limit is controlled by the singularity as $x\rightarrow 1$. We recall that ${\cal Z}(N,x)$ is a sum over partitions $p$ of ${\cal Z}(N,p,x)$. We will show that
@@
where ${\rm{Deg}}(N,p)$ is an integer we will define and ${\cal R}(x)$ is regular as $x\rightarrow 1$. We will show that the leading singularity of ${\cal Z}(N,x)$ as $x\rightarrow 1$ comes from $p=[1^{N}]$. The next-to-leading singularity come from $[1^{N-2},2]$, followed by other cycle structures involving a large number of $1$-cycles and a few small cycles. The proof that $p=[1^{N}]$ gives the leading singularity follows, while the proof that the next-to-leading singularity comes from $p=[1^{N-2},2]$ is given in Appendix B.
@@
This means that
@@
where we have factored out the terms which are singular at $x=1$.
@@ 5.1 Cycle structures and the high temperature expansion
For $k=1$,we have ${\rm{Deg}}(N,[1^{N-2},2])=N^{2}-2N+2$. Note that the coefficient of $N$ is $-2B([1^{N-2k},2^{k}])$ where $B$ is the branching number of the partition $[[1^{N-2k},2^{k}]$. The branching number of a general partition $p=[\{a_{i}^{p_{i}}\}]$ is $\sum_{i}p_{i}(a_{i}-1)$. It has an interpretation in terms of branched covers and plays a role in the Riemann-Hurwitz formula for the genus of covering surfaces in two dimensions. We will find shortly that this property, that the coefficient of $N$ in the degree is the branching number of $p$, holds for general $p$.
@@ 6 Path Integral for Tensor Models
where ${\cal D}_{\tau}=\partial_{\tau}-iA$ is the covariant derivative with the gauge field $A(\tau)=\sum_{r=1}^{s}A^{r}(\tau)$ being a Hermitian matrix in the Lie algebra of $U(N)^{s}$. A subtlety associated with this system is that though the gauge group is $U(N)^{s}$ only $SU(N)^{s}\otimes U(1)$ has a non-trivial action on the tensors i.e. there is only a single $U(1)$ gauge gield. The remaining $U(1)^{s-1}$ decouple completely from the system. The reason is rather straightforward: $U(N)$ can always be factored into $SU(N)$ and an overall $U(1)$ phase. But the $s$ overall phases act as a single phase, yielding a single diagonal $U(1)$ action. The $U(1)$ gauge fields $A_{k}$ are proportional to the identity matrix and therefore do not rotate the tensor’s index, therefore they combine into the sum $A_{1}+\cdots+A_{s}$ which is the effective $U(1)$ gauge field.
@@ 6.1 Complex-Vectors
where $P^{(1)}_{1,2d-2}(x;d)$ is a palindromic even polynomial of degree $22d-$ 2, e.g. $P^{(1)}_{1,2}(x;3)=1+x^{2}$, $P^{(1)}_{1,4}(x;4)=1+4x^{2}+x^{4}$ and $P^{(1)}_{1,6}(x;3)=1+9x^{2}+9x^{4}+x^{6}$.
@@ 6.2 Complex two-index tensors
One recognises that the invariants are built from traces of $X_{i}^{j}=\Psi_{ik}\overline{\Psi}^{kj}$ or equivalently the eigenvalues of $X_{i}^{j}$ whose square roots are the singular values of $\Psi_{ij}$ and the problem reduces to counting the invariants of a single matrix. So the invariants are the singular values of the 2-tensor, which can be thought of as a generic complex two matrix.
@@
6.3 Complex three-index tensors
@@ 9 Negative SHCs in matrix quantum mechanical models and AdS/CFT
$\displaystyle{\partial\beta_{\rm{micro}}(k)\over\partial k}=-{a\over k^{2}}$ (9.17)
@@ 9.3 Negative SHC with super-exponential growth of degeneracies in multi-matrix models
$\displaystyle Z(k,d,N)=\sum_{\begin{subarray}{c}R\vdash k\\ l(R)\leq N\end{subarray}}\sum_{\begin{subarray}{c}\Lambda\vdash k\\ l(\Lambda)\leq d\end{subarray}}C(R,R,\Lambda){\rm Dim}_{U(d)}\Lambda$ (9.34)
The sum is over Young diagrams $R$ with $k$ boxes, constrained to have no more than $N$ rows, and Young diagrams $\Lambda$ with $k$ boxes and no more than $d$ rows. $C(R,R,\Lambda)$ is the number of trivial representations of $S_{k}$ in the decomposition of the tensor product $R\otimes R\otimes\Lambda$ into irreducible representations of the diagonal $S_{k}$, i.e. the Kronecker coefficient for the triple of Young diagrams $(R,R,\Lambda)$. ${\rm Dim}_{U(d)}\Lambda$ is the dimension of the $U(d)$ representation corresponding to Young diagram $\Lambda$.
@@
$=\sum_{\begin{subarray}{c}R\vdash k\\ l(R)\leq N\end{subarray}}\sum_{p\vdash k}\sum_{q\vdash k}{1\over{\rm{Sym}}p}{1\over{\rm{Sym}}q}\chi^{R}_{p}\chi^{R}_{p}\sum_{\Lambda\vdash k}\chi^{\Lambda}_{p}\chi^{\Lambda}_{q}d^{C_{q}}$
$C_{\tau}$ is the number of cycles in the permutation $\tau$, $C_{q}$ is the number of cycles in a permutation in the conjugacy class labelled by partition $q$, equivalently the number of parts in $q$. We used the expression for the Kronecker coefficient as a sum of characters. In the last line, we have converted the sum over $\sigma$ to a sum over partitions $p$ and the sum over $\tau$ to a sum over partitions $q$. $\chi^{R}_{p}$ is the character of a permutation $\sigma_{p}$ in conjugacy class $p$ for the irreducible representation $\Lambda$. We can do the sum over $\Lambda$ using character orthogonality.
$\sum_{\Lambda\vdash k}\chi^{\Lambda}(\sigma_{p})\chi^{\Lambda}(\sigma_{q})=\delta_{p,q}~{}~{}{\rm{Sym}}~{}p$ (9.38)
@@
$\boxed{~{}~{}~{}~{}~{}~{}Z(k,d,N)=\sum_{p\vdash k}{d^{C_{p}}\over{\rm{Sym}}~{}p}\sum_{\begin{subarray}{c}R\vdash k\\ l(R)\leq N\end{subarray}}\chi^{R}_{p}\chi^{R}_{p}~{}~{}~{}~{}~{}~{}}$ (9.39)
@@
Another reason for exploring negative SHCs in multi-matrix models with $d$ scaling as $k$ is that novel scalings of $d$ in multi-matrix models are known [79] to reproduce certain aspects of the physics of tensor models, in particular the existence of limits dominated by melonic interactions. We have in section 8 that tensor models display negative SHCs at finite $N$, so it is natural to explore this feature in multi-matrix models with novel scalings.
@@
The formula (9.39) can be simplified by assuming $N\geq k$ (no assumption on $d$). The restriction $l(R)\leq N$ becomes immaterial and the formula sum over $R$ ranges over all the irreps of $S_{k}$. Using character orthgonality we then simplify to
$\displaystyle Z(k,d,N)\rightarrow Z(k,d)=\sum_{l=0}^{k}d^{l}P(k,l)$ (9.40)